can you give me the answer on this.
evaluate each expressin if a = 4,b = and c = 7
1. 3ac
2. 5b3
3. abc
4. ab/8
5. 2a-3b
6. b4/4
7. c-a
8. 20-bc
9. ab-c

Answer by bonster(299)

in order to solve these, substitute the numbers for the variables:

I'll help you with 2:
a=4
b=??? what does b =?
c=7

1) 3ac=(3)(4)(7)=(3)(28)

7)c-a=7-4=3

Don't rely on the tutors here to do your homework for you! Do your homework!